# Pushing LLMs to Their Logical Reasoning Bound: The Role of Data Reasoning Intensity

### v3

@@ A.2 Broader Impact
Our data reasoning intensity score is a pioneering metric that quantifies each example’s reasoning potential. Paired with our re-cognizing optimization framework, which uses this score to guide training, our approach reduces unnecessary computation, boosts model reasoning performance, and supports more sustainable AI practices. We believe this work will inspire new directions and offer systematic guidance for future research on unlocking the latent potential of LLM training data. Societally, our method can positively influence the efficient training of LLMs, enabling the creation of more robust and resource-efficient AI systems.
@@ A.3 Future Work
Broader model and task coverage. In future work, it would be valuable to evaluate our framework on additional architectures—beyond LLaMA-7B and Qwen2.5-7B—and across new domains such as mathematical reasoning, code generation, and multimodal understanding. This broader testing would help establish the generality and limits of data reasoning intensity and re-cognizing optimization.
@@
Compatibility with alternative data-selection methods. Our data reasoning intensity metric and re-cognizing optimization are fully compatible with other sampling strategies. In future work, we will experiment with hybrid schemes that combine our method with these complementary approaches to maximize sample efficiency and model performance and to provide a more rigorous comparison against established data-selection techniques.
@@
Pretraining-stage integration. An exciting direction is to extend reasoning-intensity guidance into the pretraining curriculum. Although pretraining multi-billion-parameter models from scratch exceeds our current resources, future work could integrate data reasoning intensity into early model training to shape core reasoning capabilities, contingent on access to sufficient compute.
Appendix B Details for the Main Experiments
B.1 Training Configuration
Compute Resources
B.2 Examples of Different Score Data
B.3 Reasoning Intensity Score Calculation Prompt
Prompt for Combinatorial Reasoning in MCQA Instructions: Please analyze the following multiple-choice question data. For each option, extract the relevant preconditions, define the deduction target, outline the deduction steps based on the provided Predicates, Constants, and Logical Expressions, and determine whether the option is correct according to the given answer index. Follow these specific rules: 1. The **Output** should be a JSON object containing only the ‘option_analysis‘ field, which is a list of analyses for each option. 2. For each option in ‘option_analysis‘, include the following fields: - **option_index**: The index of the option (starting from 0). - **option_text**: The text content of the option. - **preconditions**: A list of relevant preconditions from the Logical Expressions that pertain to the option. - **deduction_target**: The abstracted logical conclusion that the option is attempting to establish. - **deduction_steps**: A step-by-step logical deduction process from the preconditions to the deduction target. Each step should include: - **step**: The step number. - **task**: A description of what is being checked or inferred in this step. - **expression**: The logical expression used in this step, enclosed in backticks. Here, logical symbols like "implies" is represented as $\Rightarrow"$, "and" as $\land"$, "or" as $\lor"$, "not" as $\neg"$, "for all" as $\forall"$, "there exists" as $\exists"$ in LaTeX notation. - **result**: The outcome of this step, enclosed in backticks. - If a deduction step cannot proceed due to unsupported premises, indicate the failure and terminate further steps for that option. - **is_correct**: A boolean indicating whether the option is correct (true) or not (false). The values of ’Answer’ are ’0’, ’1’, ’2’, ’3’, where ’0’ represents the first option, ’1’ represents the second option, and so on. If the value of ’option_index’ is the same as the value of ’Answer’, then the value of ’is_correct’ is ’true’. 3. **Format Requirements**: - The output must strictly follow the JSON structure as shown in the example. - Ensure consistency in field naming and hierarchy. - Do not include any additional fields or information not specified in the example. 4. **Important Considerations**: - Only include preconditions that are directly relevant to the option being analyzed. - Maintain logical rigor in deduction steps, ensuring each step follows from the previous ones based on the preconditions. - Avoid including unrelated preconditions to minimize complexity and enhance clarity. Example: - **Context**: In rheumatoid arthritis, the body’s immune system misfunctions by attacking healthy cells in the joints causing the release of a hormone that in turn causes pain and swelling. This hormone is normally activated only in reaction to injury or infection. A new arthritis medication will contain a protein that inhibits the functioning of the hormone that causes pain and swelling in the joints. - **Question**: The statements above, if true, most strongly support which one of the following conclusions? - **Options**: 1. Unlike aspirin and other medications that reduce pain and swelling and that are currently available, the new medication would repair existing cell damage that had been caused by rheumatoid arthritis. 2. A patient treated with the new medication for rheumatoid arthritis could sustain a joint injury without becoming aware of it. 3. Joint diseases other than rheumatoid arthritis would not be affected by the new medication. 4. The benefits to rheumatoid arthritis sufferers of the new medication would outweigh the medication’s possible harmful side effects. - **Answer**: 1 - **Predicates**: - $\text{Attack}(x,y)$: Represents ’x’ attacking ’y’ (e.g., the immune system attacking healthy cells). - $\text{Release}(x,y)$: Represents ’x’ releasing ’y’ (e.g., the release of a hormone). - $\text{Cause}(x,y)$: Represents ’x’ causing ’y’ (e.g., the hormone causing pain and swelling). - $\text{Activate}(x,y)$: Represents ’x’ activating ’y’ (e.g., the hormone being activated by injury or infection). - $\text{Inhibit}(x,y)$: Represents ’x’ inhibiting ’y’ (e.g., the protein inhibiting the hormone). - $\text{Contain}(x,y)$: Represents ’x’ containing ’y’ (e.g., the medication containing a protein). - **Constants**: - ImmuneSystem: The body’s defense mechanism. - HealthyCells: Cells in the joints that are not diseased. - Hormone: A chemical messenger involved in causing pain and swelling. - Pain: A sensation caused by the hormone. - Swelling: A condition caused by the hormone. - Injury: A condition that normally activates the hormone. - Infection: A condition that normally activates the hormone. - ArthritisMedication: A new medication for treating arthritis. - Protein: A component of the medication that inhibits the hormone. - **Logical Expressions**: - $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ - $\text{Release}(\text{ImmuneSystem},\text{Hormone})$ - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - $\text{Activate}(\text{Injury},\text{Hormone})$ - $\text{Activate}(\text{Infection},\text{Hormone})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ Option Analysis 1. **Option Index**: 0 - **Option Text**: Unlike aspirin and other medications that reduce pain and swelling and that are currently available, the new medication would repair existing cell damage that had been caused by rheumatoid arthritis. - **Preconditions**: - $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ - $\text{Release}(\text{ImmuneSystem},\text{Hormone})$ - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - **Deduction Target**: $\text{Repair}(\text{ArthritisMedication},\text{HealthyCellsDamage})$ - **Deduction Steps**: - **Step 1**: - **Task**: Check if $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ implies $\text{Damage}(\text{HealthyCells})$. - **Expression**: $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})\Rightarrow\text{Damage}(\text{HealthyCells})$ - **Result**: Supported by context (immune system attacking healthy cells causes damage). - **Step 2**: - **Task**: Check if $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ and $\text{Inhibit}(\text{Protein},\text{Hormone})$ imply $\text{Repair}(\text{ArthritisMedication},\text{HealthyCellsDamage})$. - **Expression**: $\text{Contain}(\text{ArthritisMedication},\text{Protein})\land\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\text{Repair}(\text{ArthritisMedication},\text{HealthyCellsDamage})$ - **Result**: Not supported. The context only states that the protein inhibits the hormone, not that it repairs damage. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: $\text{Repair}(\text{ArthritisMedication},\text{HealthyCellsDamage})$ cannot be derived from the given preconditions. - **Is Correct**: False 2. **Option Index**: 1 - **Option Text**: A patient treated with the new medication for rheumatoid arthritis could sustain a joint injury without becoming aware of it. - **Preconditions**: - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Activate}(\text{Injury},\text{Hormone})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\exists\text{Patient},\text{Injury}:[\text{Sustain}(\text{Patient},\text{Injury})\land\text{Unaware}(\text{Patient},\text{Injury})]$ - **Deduction Steps**: - **Step 1**: - **Task**: Determine the effect of $\text{Inhibit}(\text{Protein},\text{Hormone})$ from the medication. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})$ - **Result**: Supported by context: The protein inhibits the hormone that causes pain and swelling. - **Step 2**: - **Task**: Analyze the implication of inhibiting the hormone on pain and swelling. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\neg\text{Cause}(\text{Hormone},\text{Pain})\land\neg\text{Cause}(\text{Hormone},\text{Swelling})$ - **Result**: Supported by context: If the hormone is inhibited, it cannot cause pain and swelling. - **Step 3**: - **Task**: Infer the patient’s awareness of injury when pain and swelling are absent. - **Expression**: $\neg\text{Cause}(\text{Hormone},\text{Pain})\land\neg\text{Cause}(\text{Hormone},\text{Swelling})\Rightarrow\text{Unaware}(\text{Patient},\text{Injury})$ - **Result**: Supported by context: Without pain and swelling, the patient may not be aware of sustaining an injury. - **Step 4**: - **Task**: Combine the above implications to conclude the deduction target. - **Expression**: $\exists\text{Patient},\text{Injury}:[\text{Sustain}(\text{Patient},\text{Injury})\land\text{Unaware}(\text{Patient},\text{Injury})]$ - **Result**: Deduction is valid based on the inhibited hormone preventing awareness of injury. - **Is Correct**: True 3. **Option Index**: 2 - **Option Text**: Joint diseases other than rheumatoid arthritis would not be affected by the new medication. - **Preconditions**: - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\forall x[\text{JointDisease}(x)\land x\neq\text{RheumatoidArthritis}\Rightarrow\neg\text{Affect}(\text{ArthritisMedication},x)]$ - **Deduction Steps**: - **Step 1**: - **Task**: Check if the context provides information about other joint diseases. - **Expression**: $\text{JointDisease}(x)\land x\neq\text{RheumatoidArthritis}$ - **Result**: Not supported. The context only discusses rheumatoid arthritis. - **Step 2**: - **Task**: Determine if there is any implication that the medication specifically targets rheumatoid arthritis. - **Expression**: $\text{Contain}(\text{ArthritisMedication},\text{Protein})\Rightarrow\text{SpecificEffect}(\text{RheumatoidArthritis})$ - **Result**: Not supported. The context does not specify that the protein exclusively affects rheumatoid arthritis. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: Cannot conclude that the medication does not affect other joint diseases. - **Is Correct**: False 4. **Option Index**: 3 - **Option Text**: The benefits to rheumatoid arthritis sufferers of the new medication would outweigh the medication’s possible harmful side effects. - **Preconditions**: - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\text{Benefit}(\text{ArthritisMedication})>\text{HarmfulSideEffect}(\text{ArthritisMedication})$ - **Deduction Steps**: - **Step 1**: - **Task**: Identify the benefits of the medication based on inhibiting the hormone. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\text{Reduce}(\text{Pain})\land\text{Reduce}(\text{Swelling})$ - **Result**: Supported by context: The protein inhibits the hormone, which causes pain and swelling. - **Step 2**: - **Task**: Determine if the context provides information about harmful side effects. - **Expression**: $\text{HarmfulSideEffect}(\text{ArthritisMedication})$ - **Result**: Not supported. The context does not mention any side effects of the medication. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: Cannot compare benefits and harmful side effects due to lack of information on side effects. - **Is Correct**: False Tips for Option Analysis - **Preconditions**: Only include logical expressions that are directly relevant to the option being analyzed. Avoid listing all possible preconditions. - **Deduction Steps**: Ensure each step logically follows from the previous one based on the preconditions. If a step cannot be completed due to insufficient support from the preconditions, indicate the failure and stop further deductions for that option. - **is_correct**: This field should be true only for the option that matches the answer index. All other options should be false. - **Format Consistency**: Maintain the same JSON structure and field naming conventions across all options to ensure uniformity and ease of data extraction. - **Logical Accuracy**: Ensure that all logical expressions and deductions accurately reflect the relationships defined by the predicates and constants. Your Task Analyze the following Input data and generate the option_analysis section as per the example above. Replace the xxx placeholders in the example with actual data derived from the input. Input Data input_data_here Please generate the option_analysis section based on the above input data.
Appendix C DRI Calculation Process and Re-Cognizing Optimization Algorithm
Algorithm 1 illustrates the calculation process of our DRI score, and Algorithm 2 demonstrates the implementation procedure of our Re-Cognizing Optimization strategy.
Algorithm 1 Data Reasoning Intensity Calculation
@@ Appendix C DRI Calculation Process and Re-Cognizing Optimization Algorithm
3: Extract logical expressions $\mathcal{E}$ and compute their nesting depths.
4: Identify predicates $\mathcal{P}$ and constants $\mathcal{C}$ in $c$.
@@
8:  Extract preconditions $\mathcal{R}_{l}$ and reasoning steps $\mathcal{S}_{l}$.
@@
10:  $S_{\text{opt}}^{(l)}\;=\;|\mathcal{R}_{l}|\cdot\bar{D}_{l}^{2}\;+\;\sum_{k=1}^{T_{l}}{\bigl(1+\#\text{Operations}_{l,k}\bigr)\;D_{l,k}^{2}}$.
@@
1:Input: dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, model $M$, epochs $T$.
@@
5:   Uniformly shuffle $\mathcal{D}$ for initial exploration (Phase I: Model Cognition Reshaping).
@@
7:   Sort $\mathcal{D}$ by descending $S_{i}$ to emphasize high-intensity samples (Phase II: Cognitive Reasoning Enhancement).
@@
9:  for each batch $B$ drawn sequentially from $\mathcal{D}$ do

### v4

@@ A.2 Impact
Our SLKD score is a pioneering metric that quantifies each example’s reasoning potential. Paired with our re-cognizing optimization framework, which uses this score to guide training, our approach reduces unnecessary computation, boosts model reasoning performance, and supports more sustainable AI practices. We believe this work will inspire new directions and offer systematic guidance for future research on unlocking the latent potential of LLM training data. Societally, our method can positively influence the efficient training of LLMs, enabling the creation of more robust and resource-efficient AI systems.
@@ A.3 Future Work
Broader model and task coverage. In future work, it would be valuable to evaluate our framework on additional architectures—beyond LLaMA-7B and Qwen2.5-7B—and across new domains such as mathematical reasoning, code generation, and multimodal understanding. This broader testing would help establish the generality and limits of Structured Logical Knowledge Density and re-cognizing optimization.
@@
Compatibility with alternative data-selection methods. Our Structured Logical Knowledge Density metric and re-cognizing optimization are fully compatible with other sampling strategies. In future work, we will experiment with hybrid schemes that combine our method with these complementary approaches to maximize sample efficiency and model performance and to provide a more rigorous comparison against established data-selection techniques.
@@
Pretraining-stage integration. An exciting direction is to extend reasoning-intensity guidance into the pretraining curriculum. Although pretraining multi-billion-parameter models from scratch exceeds our current resources, future work could integrate SLKD into early model training to shape core reasoning capabilities, contingent on access to sufficient compute.
Algorithm 1 SLKD Calculation
@@
3: Extract logical expressions E and compute their nesting depths.
4: Identify predicates P{P} and constants C{C} in $c$.
@@
8:  Extract preconditions ${R}_{l}$ and reasoning steps ${S}_{l}$.
@@
10:  $S_{\text{opt}}^{(l)}\;=\;|{R}_{l}|\cdot\bar{D}_{l}^{2}\;+\;\sum_{k=1}^{T_{l}}{\bigl(1+\#\text{Operations}_{l,k}\bigr)\;D_{l,k}^{2}}$.
@@
1:Input: dataset ${D}=\{x_{i}\}_{i=1}^{N}$, model $M$, epochs $T$.
@@
5:   Uniformly shuffle D{D} for initial exploration (Phase I: Model Cognition Reshaping).
@@
7:   Sort D{D} by descending $S_{i}$ to emphasize high-intensity samples (Phase II: Cognitive Reasoning Enhancement).
@@
9:  for each batch $B$ drawn sequentially from D{D} do